one type of cattle feed is 4% protein while a second type is 13% protein. how many pounds of each type must be mixed to make 45 pounds of mixture that is 10% protein?
Answer by nerdybill(7384)

one type of cattle feed is 4% protein while a second type is 13% protein. how many pounds of each type must be mixed to make 45 pounds of mixture that is 10% protein?
.
Let x = amount (lbs) of 4% protein
then
45-x = amount (lbs) of 13% proteen
.
.04x + .13(45-x) = .10(45)
.04x + 5.85-.13x) = 4.5
5.85-.09x = 4.5
5.85 = .09x+4.5
1.35 = .09x
15 pounds = x (amount of 4% protein feed)
.
13% protein feed:
45-x = 45-15 = 30 pounds